OpenStudy (anonymous):

Simplify these! 8^2+11 (-2)^4

OpenStudy (anonymous):

8^2+11= 64 + 11= 75

OpenStudy (anonymous):

64+11=75 (-2)^4=16

OpenStudy (anonymous):

(-2)^4= -2 *-2 = 4 negative times negative is positive